one stick is 9cm longer than the other. 2/5 of the long stick is equal to 1/2 of the shorter stick. find the length of the longer stick.
Answer by JulietG(1812)

Let S be the short stick
Let L be the long stick
Here's what we're told:
L = 9 + S
2/5L = 1/2S
.
Substitute the value for L
2/5(9+S) = 1/2S
Multiply each side by 5 to get rid of that first fraction
2(9+S) = 1/2S * 5
Distribute and multiply
18 + 2S = 5/2 * S
Subtract 2S from each side (which is 4/2S)
18 = 1/2S
Multiply each side by 2 to get rid of the fraction
36 = S
If the short stick is 36, then the longer stick is 36+9=45 <<--
.
Let's check it.
"2/5 of the long stick is equal to 1/2 of the shorter stick"
2/5 * 45 = 90/5 = 18
1/2 * 36 = 36/2 = 18
They are equal.
Success!